# Towards Robust Spatio-Temporal Auto-Regressive Prediction: Adams-Bashforth Time Integration with Adaptive Multi-Step Rollout

### v1

@@ 2.2.1 Motivation for multi-step rollout
2.2.2 Background on conventional multi-step rollout
@@
Specifically, given input snapshots $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, the AR model first predicts $\hat{\mathbf{u}}(t+1)$. The input history is then updated to $\hat{\mathbf{u}}(t+1),\ldots,\mathbf{u}(t-(N-2))$, allowing the model to predict $\hat{\mathbf{u}}(t+2)$. This iterative process continues for $M$ steps, producing predictions from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$ based on the initial input history $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, without any backpropagation during inference. The predicted future snapshots, from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$, are then used to compute the loss: from $\mathcal{L}_{1}$ to $\mathcal{L}_{M}$. In the training process, this method minimizes a multi-step loss function, which aggregates losses across all prediction steps. By explicitly optimizing across multiple steps, the model enhances its resilience to compounding errors during inference. Consequently, the multi-step loss function is defined as a weighted sum of individual step losses, expressed as:
@@
where $w_{i}$ represents the weight assigned to the loss at each future time step $t+i$. Notably, when $M=1$, this corresponds to the conventional loss without employing the multi-step rollout technique. Backpropagation with respect to $\mathcal{L}_{\text{multi-step}}$ enables the model to optimize for long-term predictions by accounting for all loss components, from $\mathcal{L}_{1}$ to $\mathcal{L}_{M}$.
Figure 1: Overview of multi-step rollout: the model predicts $M$ future steps during training, with the total loss computed as a weighted sum of individual prediction losses at each step. In this study, the adaptive weighting schemes for dynamically determining weights ($w_{i}$) will be newly proposed in Section 2.2.3.
@@
The conventional multi-step rollout considers all future $M$ steps simultaneously during training. This can complicate optimization, as the model must account for every snapshot in the long-term future prediction process. As a result, the model might struggle to balance short-term and long-term accuracy, particularly in early stages of training when long-term predictions are highly uncertain. In summary, the simultaneous consideration of all $M$ steps can lead to poor convergence, as errors from long-term predictions dominate, making it harder for the model to achieve stable optimization.
In conclusion, while the conventional multi-step rollout method proposed by Wu et al. [21] is effective for capturing sequential dependencies in time-series data, it presents challenges. Specifically, it requires careful weight tuning and faces the difficulty of optimizing the model simultaneously for all future steps, which can impede convergence and reduce robustness.
@@ 2.2.3 Adaptive multi-step rollout
The first approach seeks to mitigate the need for manually selecting weights $\lambda_{i}$ by automatically assigning weights based on the magnitude of the prediction errors at each time step. Specifically, the weights are computed as the normalized mean squared error (MSE) for each prediction step:
@@
$\text{MSE}_{i}=\left|\hat{\mathbf{u}}(t+i)-\mathbf{u}(t+i)\right|^{2},\quad i=1,\dots,M$
@@
6:     Compute MSE for each step: $\text{MSE}_{i}=|\hat{\mathbf{u}}(t+i)-\mathbf{u}(t+i)|^{2}$ for $i=1,\dots,M$
@@ 3.2 Model architecture
The Graph U-Net architecture employed in this study is designed with a focus on computational efficiency and simplicity, comprising only 1,177 trainable parameters (Fig. 3). The encoder part of the model consists of four graph convolutional network (GCN) layers with channel dimensions decreasing from 20 to 1 in the sequence 20, 15, 10, 5, and 1. A pooling ratio of 0.6 is applied using the gPool layer [22] after each GCN layer in the encoder, as this ratio was found to be optimal in our previous work [17]. This pooling operation reduces the number of nodes by 60% at each step, effectively capturing multi-scale features while maintaining computational tractability.
@@ 3.3 Training procedure
The model is trained for 5,000 epochs using the Adam optimizer with an initial learning rate of $10^{-3}$. To ensure robustness and capture variability, each training process is repeated three times under identical settings, with the mean performance metrics reported as the final experimental results. During training, the MSE between predicted and ground-truth snapshots serves as the loss function.
@@ 3.4 Evaluation and rollout process
Considering the small model capacity (1,177 parameters) and the limited size of the training dataset (50 snapshots), the rollout process presents a challenging scenario for the model. The extended prediction horizon of 350 snapshots is substantially longer than the training sequence and encompasses multiple periods of vortex shedding. This harsh evaluation setting is intended to test the model’s ability to generalize temporal dynamics and maintain stability over long-term predictions.
@@ 3.5 Underlying reason for harsh conditions
4 Application of time schemes into auto-regressive GNNs
In this section, we present the results of applying the four different time integration schemes, along with the conventional direct prediction approach, as described in Section 2.1.2. The objective is to assess their own impact on the long-term prediction accuracy and stability of the Graph U-Net model without yet incorporating the multi-step rollout technique (which will be addressed in later sections).
4.1 Predictive error comparison
@@
First-order central 65.024 1879
Second-order central 0.713 2003
@@
As observed in Table 1, the direct prediction (conventional approach), forward Euler, and Adams-Euler schemes yield relatively low MSE values, seemingly having better prediction accuracy. In contrast, the first-order central difference exhibits significantly higher errors. The training time on the NVIDIA GeForce RTX 3090 24GB GPU is also summarized (all of the experiments conducted in this study is based on the same GPU), indicating that the use of different time integration schemes does not significantly affect the time required to train the model.
Among various time integration approaches, Fig. 4 illustrates the predicted snapshots when using the first-order central difference scheme. Four consecutive snapshots after 100 rollouts are shown, revealing a highly oscillatory behavior: snapshots after 100 (Fig. 4(a)) and 102 (Fig. 4(c)) rollouts appear similar, while those after 101 (Fig. 4(b)) and 103 (Fig. 4(d)) exhibit similar patterns. This alternating pattern persists beyond 100 rollouts (not shown here), reflecting a fundamental limitation of the first-order central difference method. Specifically, the next time step, $\mathbf{u}(t+\Delta t)$, depends heavily on $\mathbf{u}(t-\Delta t)$, bypassing the intermediate state at $\mathbf{u}(t)$ (as described in Eq. 6). These results align with established findings that the central difference schemes are generally unsuitable for time integration due to their inherent instability, resulting in oscillatory and non-physical solutions.
@@
Figure 4: Predicted snapshots using the first-order central difference scheme after (a) 100, (b) 101, (c) 102, and (d) 103 rollout steps. Snapshots after 100 and 102 rollouts display similar patterns, while those after 101 and 103 show similar but alternate patterns.
@@ 4.2 Need for additional techniques
The results indicate that, despite the relatively low MSE values for some schemes, the Graph U-Net models struggle to maintain accurate long-term predictions over extended rollout horizons. The accumulation of errors becomes apparent even within the first 100 rollout steps, well short of the 350-step prediction horizon that is the main objective of this study. The high errors observed in the first-order and second-order central difference schemes suggest that these methods are less robust for this application, potentially due to their reliance on central difference approximations, which may amplify errors in an AR prediction context due to their oscillatory behavior found in Fig. 4.
@@
These findings underscore the limitations of directly applying traditional finite difference schemes within GNN architectures for time integration, revealing insufficient performance over extended prediction horizons. This highlights the need for enhanced techniques to improve stability and accuracy, such as integrating multi-step rollout strategies to address the observed shortcomings in current methods. In the subsequent sections, we will introduce and evaluate the impact of integrating multi-step rollout techniques with the Graph U-Net models. By combining these approaches with appropriate time integration schemes, we aim to enhance the long-term prediction performance and achieve more reliable forecasts over extended time horizons.
@@
5 Extension of time integration schemes into multi-step rollout scenario
@@
In this section, we extend the previously investigated time integration schemes by incorporating multi-step rollout techniques with different values of $M$, where $M$ represents the number of future snapshots considered during training. Specifically, we examine $M=1$, $2$, $4$, and $8$. When $M=1$, it means that the multi-step rollout is not applied, corresponding to the vanilla AR model. By integrating multi-step rollout, we aim to assess whether training the model to predict multiple future steps can enhance long-term prediction accuracy and stability. Note that the weighting strategy suggested by Wu et al. [21] is adopted herein, where the loss weights are set to $w_{1}=1$ for the first future step and $w_{i}=0.1$ for the subsequent steps in the multi-step rollout (see Eq. 12).
5.1 Performance evaluation with multi-step rollout
@@
We evaluate all the time integration schemes with varying values of $M$ and the MSE results for each combination are presented in Table 2. Here, we can see that increasing the multi-step rollout length $M$ introduces significant computational overhead, as evidenced by the increased training times from 1,914s ($M=1$) to 3,131s ($M=8$) shown in Table 2. However, this increased computational cost does not necessarily lead to improved performance. The results reveal a clear distinction between different time integration schemes: While the Adams-Bashforth with forward Euler (Adams-Euler) scheme demonstrates consistently robust performance across all values of $M$ (MSE improving from 0.139 at $M=1$ to 0.070 at $M=4$), other schemes exhibit significant instability. Notably, the direct prediction scheme breaks down completely with NaN errors at $M=4$ and $M=8$, while the forward Euler method shows instability at $M=2$ (MSE of 581.945) and $M=8$ (NaN). Similarly, both first-order and second-order central difference schemes display extremely high MSE values across increasing $M$.
@@
These results highlight a crucial finding: naively increasing $M$ imposes excessive training constraints on the model, particularly challenging given its limited capacity of only 1,177 parameters. While longer rollout lengths theoretically allow the model to learn longer-term dependencies, they also require the model to maintain stability across more time steps during training. Most time integration schemes struggle with this requirement, leading to degraded performance or complete breakdown. More importantly, these results demonstrate that the selection of an appropriate time integration scheme plays a more critical role in achieving stable AR predictions than the choice of $M$. This is evident from Table 2, where even with optimal $M$ values, most time integration schemes fail to achieve stable predictions, while the Adams-Euler scheme maintains robust performance across all $M$ values.
@@
The Adams-Euler scheme, introduced for the first time in this paper for AR prediction, stands out by maintaining stability even at higher $M$ values. This superior performance can be attributed to its ability to leverage information from multiple previous time steps—since three successive snapshots are explicitly considered for the prediction of the next snapshot as in Eq. 9—in a mathematically principled way, enabling stable predictions even with limited model capacity.
To illustrate the improved performance achieved with the Adams-Euler scheme with $M=4$ multi-step rollout (best case in Table 2), we present visual comparisons of the predicted flow fields after 200 and 300 rollout steps: Fig. 6 shows the ground truth and the model’s predictions at these rollout steps. As seen in Figs. 6(a) and 6(b), the prediction at snapshot $t+200$ shows reasonable agreement with the ground truth, capturing the overall flow patterns and vortex shedding behavior at specific phase. This indicates a significant improvement compared to the results in Section 4, where accurate predictions were not achievable even in the 100 rollout predictions.
@@
However, at snapshot $t+300$ (Figs. 6(c) and 6(d)), the MSE increases from 0.091 at snapshot $t+200$ to 0.105, resulting in a noticeable divergence between the prediction and the ground truth. The accuracy of the flow structures diminishes, with blurred vortex shedding trail. This indicates that while the incorporation of multi-step rollout and the Adams-Euler scheme significantly enhances long-term prediction performance, there remains room for improvement, particularly for prediction horizons extending beyond $t+200$.
@@ 6 Application of proposed adaptive multi-step rollout
Performance of the three adaptive weighting strategies across different time integration schemes, direct prediction, forward Euler, and Adams-Euler, is assessed. Unlike previous sections, the MSE results from this section are calculated based on $x$-velocity data from seven probe points, as shown in Fig. 7. These results are summarized in Table 3. For the direct prediction case, while the vanilla approach with fixed weights (MSE of 0.011) outperforms its adaptive variants, it still falls short compared to the best results achieved with forward Euler and Adams-Euler schemes. Notably, both our proposed approaches (AW2 and AW3, which are adaptive weightings with a learnable parameter) demonstrate strong performance with the forward Euler scheme (MSE of 0.007 and 0.010 respectively), which is a widely adopted approach in GNN-based AR prediction. However, the most remarkable performance is achieved by combining the Adams-Euler scheme with AW3, which emphasizes only the first and last loss components during training. This combination yields a significantly lower MSE (0.002) compared to all other methods, establishing it as the optimal approach for long-term prediction in our study. A comparison of the required training time reveals negligible differences across the various weighting approaches. This indicates that the proposed adaptive weighting methods enhance rollout performance without incurring additional computational costs.
Figure 7: Locations of seven probe points examined to gain a more intuitive understanding of vortex shedding prediction in terms of $x$-velocity.
@@
Forward Euler 0.019 0.023 0.007 0.010
Adams-Euler 0.018 0.017 0.010 0.002
@@
  • Stabilization phase (before 1000 epochs) During the initial training phase, the Graph U-Net stabilizes, with minimal differences observed between the weights of the four loss terms. This indicates that the model is primarily learning basic temporal patterns, while the adaptive weighting mechanism has not yet significantly shifted its focus toward specific time steps.
  • Adaptation phase (after 1000 epochs) Beyond epoch 1000, the weight assigned to the 4th loss term increases relative to the others. The weights for the first three loss terms do not exhibit clear trends or significant differences among themselves.
Figure 8: Evolution of adaptive weights for each loss term over epochs in AW2 with the Adams-Euler scheme. Adaptive weights ($w_{i}$) and effective parameter ($k_{eff}$) are shown in each y-axis.
@@
To intuitively and quantitatively evaluate the impact of the improvements achieved by AW3 on capturing vortex shedding behavior, we analyze the temporal evolution of the $x$-velocity at seven probe points positioned downstream of the cylinder (Fig. 7). Based on these probe points, we compare the predictions of the two best-performing models trained so far (Table 3). Model A: the one with forward Euler scheme and AW2 (adaptive weighting with learnable parameter), resulting in an MSE of 0.007. Model B: the other with Adams-Euler scheme with AW3 (first and last loss terms only), resulting in an MSE of 0.002.

### v2

@@ 2.2.1 Motivation for multi-step rollout
2.2.2 Conventional multi-step rollout
@@
Specifically, given input snapshots $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, the AR framework first predicts $\hat{\mathbf{u}}(t+1)$. The input history is then updated to $\hat{\mathbf{u}}(t+1),\ldots,\mathbf{u}(t-(N-2))$, allowing to predict $\hat{\mathbf{u}}(t+2)$. This iterative process continues for $M$ steps, producing predictions from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$ based on the initial input history $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, without any backpropagation during inference. The predicted future snapshots, from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$, are then used to compute the loss: from $\mathcal{L}_{1}$ to $\mathcal{L}_{M}$. In the training process, this method minimizes a multi-step loss function, which aggregates losses across all prediction steps. By explicitly optimizing across multiple steps, the model enhances its resilience to compounding errors during inference. Consequently, the multi-step loss function is defined as a weighted sum of individual step losses, expressed as:
@@
where $w_{i}$ represents the weight assigned to the loss at each future time step $t+i$. Notably, when $M=1$, this corresponds to the conventional loss without applying the multi-step rollout technique. Backpropagation with respect to $\mathcal{L}_{\text{multi-step}}$ enables the model to optimize for long-term predictions by accounting for all loss components, from $\mathcal{L}_{1}$ to $\mathcal{L}_{M}$.
Figure 2: Overview of multi-step rollout: the model predicts $M$ future steps during training, with the total loss computed as a weighted sum of individual prediction losses at each step. In this study, the adaptive weighting schemes for dynamically determining weights ($w_{i}$) are newly proposed in Section 2.2.3.
@@
The conventional multi-step rollout considers all future $M$ steps simultaneously during training. This can complicate optimization, as the model should account for every snapshot in the long-term future prediction process. As a result, the model might struggle to balance short-term and long-term accuracy, particularly in early stages of training when long-term predictions are highly uncertain. In summary, the simultaneous consideration of all $M$ steps can lead to poor convergence, as errors from long-term predictions dominate, making it harder for the model to achieve stable optimization.
@@ 2.2.3 Adaptive multi-step rollout
The first approach seeks to mitigate the need for manually selecting weights $w_{i}$ by automatically assigning weights based on the magnitude of the prediction errors at each time step. Specifically, the weights are computed as the normalized mean squared error (MSE) for each prediction step:
@@
$\text{MSE}_{i}=MSE(\hat{\mathbf{u}}(t+i),\mathbf{u}(t+i)),\quad i=1,\dots,M$
@@
6:     Compute MSE for each step: $\text{MSE}_{i}=MSE(\hat{\mathbf{u}}(t+i),\mathbf{u}(t+i))$ for $i=1,\dots,M$
@@ 3.2 Model architecture
The Graph U-Net architecture employed in this study is designed with a focus on computational efficiency and simplicity, comprising only 1,177 trainable parameters (Fig. 4). The encoder part of the model consists of four graph convolutional network (GCN) layers [31] with channel dimensions decreasing from 20 to 1 in the sequence 20, 15, 10, 5, and 1. A pooling ratio of 0.6 is applied using the gPool layer [30] after each GCN layer in the encoder, as this ratio was found to be optimal in our previous work [24]. This pooling operation reduces the number of nodes by 60% at each step, effectively capturing multi-scale features while maintaining computational tractability.
@@ 3.3 Training procedure
The model is trained for 5,000 epochs using the Adam optimizer with an initial learning rate of $10^{-3}$. To ensure robustness and capture variability, each training process is repeated three times under identical settings, with the mean performance metrics presented as experimental results. During training, the MSE between predicted and ground-truth snapshots serves as the loss function.
@@ 3.5 Underlying reason for harsh conditions
4 Application of time integration schemes into auto-regressive GNNs
In this section, we present the results of applying the four different time integration schemes, along with the conventional direct prediction approach, as described in Section 2.1.2. The objective is to assess their own impact on the long-term prediction accuracy and stability of the Graph U-Net model before incorporating the multi-step rollout technique (which will be addressed in later sections).
4.1 Predictive performance comparison
@@
Second-order central 65.024 1879
@@
As observed in Table 1, the direct prediction (conventional approach), forward Euler, and Adams-Euler schemes yield relatively low MSE values, seemingly having better prediction accuracy. In contrast, the second-order central difference exhibits significantly higher errors. The training time on the NVIDIA GeForce RTX 3090 24GB GPU is also summarized (all of the experiments conducted in this study is based on the same GPU), indicating that the use of different time integration schemes does not significantly affect the time required to train the model.
Among various time integration approaches, Fig. 5 illustrates the predicted snapshots when using the second-order central difference scheme. Four consecutive snapshots after 100 rollouts are shown, revealing a highly oscillatory behavior: snapshots after 100 (Fig. 5(a)) and 102 (Fig. 5(c)) rollouts appear similar, while those after 101 (Fig. 5(b)) and 103 (Fig. 5(d)) exhibit similar patterns. This alternating pattern persists beyond 100 rollouts (not shown here), reflecting a fundamental limitation of the central difference method. Specifically, the next time step, $\mathbf{u}(t+\Delta t)$, depends heavily on $\mathbf{u}(t-\Delta t)$, bypassing the intermediate state at $\mathbf{u}(t)$ (as described in Eq. 6). These results align with established findings that the central difference schemes are generally unsuitable for time integration due to their inherent instability, resulting in oscillatory and non-physical solutions.
@@
Figure 5: Predicted snapshots using the second-order central difference scheme after (a) 100, (b) 101, (c) 102, and (d) 103 rollout steps. Snapshots after 100 and 102 rollouts display similar patterns, while those after 101 and 103 show similar but alternate patterns.
@@ 4.2 Need for additional techniques
The results in Section 4.1 indicate that, despite the relatively low MSE values for some schemes, the Graph U-Net models struggle to maintain accurate long-term predictions over extended rollout horizons. The accumulation of errors becomes apparent even within the first 100 rollout steps, well short of the 350-step prediction horizon that is the main objective of this study. These findings underscore the limitations of directly applying traditional finite difference schemes within AR architectures for time integration, revealing insufficient performance over extended prediction horizons. This highlights the need for enhanced techniques to improve stability and accuracy, such as integrating multi-step rollout strategies to address the observed shortcomings in current methods. In the subsequent sections, we will introduce and evaluate the impact of integrating multi-step rollout techniques with the Graph U-Net models. By combining these approaches with appropriate time integration schemes, we aim to enhance the long-term prediction performance and achieve more reliable forecasts over extended time horizons.
@@
5 Extension of time integration schemes into conventional multi-step rollout scenario
@@
In this section, we extend the previously investigated time integration schemes by incorporating conventional multi-step rollout techniques with different values of $M$, where $M$ represents the number of future snapshots considered during training. Specifically, we examine $M=1$, $2$, $4$, and $8$. When $M=1$, it means that the multi-step rollout is not applied, corresponding to the vanilla AR model. By integrating multi-step rollout, we aim to assess whether training the model to predict multiple future steps can enhance long-term prediction accuracy and stability. Note that the conventional weighting strategy suggested by Wu et al. [29] is adopted herein, where the loss weights are set to $w_{1}=1$ for the first future step and $w_{i}=0.1$ for the subsequent steps in the multi-step rollout (see Eq. 10).
5.1 Performance evaluation with conventional multi-step rollout
@@
We evaluate all the time integration schemes with varying values of $M$ and the MSE results for each combination are presented in Table 2. Here, we can see that increasing the multi-step rollout length $M$ introduces significant computational overhead, as evidenced by the increased training times from 1,914s ($M=1$) to 3,131s ($M=8$) shown in Table 2. However, this increased computational cost does not necessarily lead to improved performance. The results reveal a clear distinction between different time integration schemes: while the Adams-Euler scheme demonstrates consistently robust performance across all values of $M$ (MSE improving from 0.139 at $M=1$ to 0.070 at $M=4$), other schemes exhibit significant instability. Notably, the direct prediction scheme breaks down completely with NaN errors at $M=4$ and $M=8$, while the forward Euler method shows instability at $M=2$ (MSE of 581.945) and $M=8$ (NaN). Similarly, second-order central difference scheme displays extremely high MSE values across increasing $M$.
@@
These results reveal two key insights: First, naively increasing $M$ imposes excessive training constraints on the lightweight model (1,177 parameters), often leading to performance degradation or complete breakdown despite the theoretical benefit of learning longer-term dependencies. Second, the time integration scheme proves more critical than the choice of $M$—while most methods fail even with optimal $M$ values, the Adams-Euler scheme maintains robust performance across all $M$ values, demonstrating its inherent stability for long-term AR prediction.
@@
The Adams-Euler-based AR scheme, introduced for the first time in this paper, stands out by maintaining stability even at higher $M$ values. This superior performance can be attributed to its ability to leverage information from multiple previous time steps—since three successive snapshots, $\mathbf{u}(t-\Delta t)$, $\mathbf{u}(t)$, and $\mathbf{u}(t+\Delta t)$, are explicitly considered for the prediction of the next snapshot as in Eq. 7—in a mathematically principled way, enabling stable predictions even with limited model capacity.
To illustrate the improved performance achieved with the Adams-Euler scheme with $M=4$ multi-step rollout (best case in Table 2), we present visual comparisons of the predicted flow fields after 200 and 300 rollout steps: Fig. 7 shows the ground truth and the model’s predictions at these rollout steps. As seen in Figs. 7(a) and 7(b), the prediction at snapshot $t+200$ shows reasonable agreement with the ground truth, capturing the overall flow patterns and vortex shedding behavior at specific phase. This represents a substantial improvement over the results in Fig. 6, where accurate predictions were unattainable even at 100 rollout steps, clearly demonstrating the effectiveness of combining conventional multi-step rollout with the Adams-Euler scheme.
@@
However, at snapshot $t+300$ (Figs. 7(c) and 7(d)), the MSE increases from 0.091 at snapshot $t+200$ to 0.105, resulting in a noticeable divergence between the prediction and the ground truth. The accuracy of the flow structures diminishes, with blurred vortex shedding trail. This indicates that while the incorporation of multi-step rollout and the Adams-Euler scheme significantly enhances long-term prediction performance, there remains room for improvement, particularly for prediction horizons extending beyond $t+200$.
@@ 6 Application of proposed adaptive multi-step rollout
Performance of the three adaptive weighting strategies across different time integration schemes—direct prediction, forward Euler, and Adams-Euler—is assessed using a more focused evaluation approach. To better evaluate the models’ ability to capture complex wake oscillatory vortex behavior, MSE results from this section are calculated based on $x$-velocity data from seven strategic probe points in the wake region (shown in Fig. 8), rather than across the entire field. This targeted analysis provides a more precise assessment of how well each method captures the critical vortex shedding dynamics and the results are summarized in Table 3. For the direct prediction case, while the vanilla approach with fixed weights (MSE of 0.011) outperforms its adaptive variants, it still falls short compared to the best results achieved with forward Euler and Adams-Euler schemes. Notably, both our proposed approaches (AW2 and AW3, which are adaptive weightings with a learnable parameter) demonstrate strong performance with the forward Euler scheme (MSE of 0.007 and 0.010 respectively), which is a widely adopted approach in GNN-based AR prediction. However, the most remarkable performance is achieved by combining the Adams-Euler scheme with AW3, which emphasizes only the first and last loss components during training. This combination yields a significantly lower MSE (0.002) compared to all other methods, establishing it as the optimal approach for long-term prediction in our study. A comparison of the required training time reveals negligible differences across the various weighting approaches: it indicates that the proposed adaptive weighting methods enhance rollout performance without incurring additional computational costs.
Figure 8: Locations of seven probe points examined to gain a more intuitive and quantitative understanding of vortex shedding prediction performance in terms of $x$-velocity.
@@
Forward Euler 0.019 0.023 0.007* 0.010
Adams-Euler 0.018 0.017 0.010 0.002**
@@
  • Stabilization phase (before 1000 epochs): during the initial training phase, the Graph U-Net stabilizes, with minimal differences observed between the weights of the four loss terms. This indicates that the model is primarily learning basic temporal patterns, while the adaptive weighting mechanism has not yet significantly shifted its focus toward specific time steps.
  • Adaptation phase (after 1000 epochs): beyond epoch 1000, the weight assigned to the 4th loss term increases relative to the others. The weights for the first three loss terms do not exhibit clear trends or significant differences among themselves.
Figure 9: Evolution of adaptive weights for each loss term over epochs in AW2 with the Adams-Euler scheme (multi-step rollout with $M=4$). Adaptive weights ($w_{i}$) and effective parameter ($k_{eff}$) are shown in each y-axis.
@@
To intuitively and quantitatively evaluate the impact of the improvements achieved by AW3 on capturing vortex shedding behavior, we analyze the temporal evolution of the $x$-velocity at seven probe points positioned downstream of the cylinder (Fig. 8). Based on these probe points, we compare the predictions of the two best-performing models trained so far (two models with footnotes in Table 3). Model A: the one with forward Euler scheme and AW2 (adaptive weighting with learnable parameter), resulting in an MSE of 0.007. Model B: the other with Adams-Euler scheme with AW3 (first and last loss terms only), resulting in an MSE of 0.002.